I know correct answer is c. However what type of directors are A and B (meta directors ) ??? Please explain conceptually . Thank you!
Answer #1

Yes. A & B are strongly deactivating groups. Hence meta directing groups.

Some  meta directing groups. are: -NO2, -NR3+, -NH3+, -SO2-OH, -C\equivN, -CF3, C=O-Cl

Explanation:

Electron Donating Groups / activating groups direct ortho / para

Electron Withdrawing Groups / deactivating groups direct meta

except halogens (-X) which are deactivating BUT direct ortho / para.

EWGs normally attached to \pi bond or have formal +ve charge